# Senior Authors

### v1

@@ Abstract
Multimodal agents for computer use rely exclusively on primitive actions (click, type, scroll) that require accurate visual grounding and lengthy execution chains, leading to cascading failures and performance bottlenecks. While other agents leverage rich programmatic interfaces (APIs, MCP servers, tools), computer-use agents (CUAs) remain isolated from these capabilities. We present UltraCUA, a foundation model that bridges this gap through hybrid action—seamlessly integrating GUI primitives with high-level programmatic tool calls. To achieve this, our approach comprises four key components: (1) an automated pipeline that scales programmatic tools from software documentation, open-source repositories, and code generation; (2) a synthetic data engine producing 17,000+ verifiable tasks spanning real-world computer-use scenarios; (3) a large-scale high-quality hybrid action trajectory collection with both low-level GUI actions and high-level programmatic tool calls; and (4) a two-stage training pipeline combining supervised fine-tuning with online reinforcement learning, enabling strategic alternation between low-level and high-level actions. Experiments with our 7B and 32B models demonstrate substantial improvements over state-of-the-art agents. On OSWorld, UltraCUA models achieve an average 22% relative improvement over base models, while being 11% faster in terms of steps. Out-of-domain evaluation on WindowsAgentArena shows our model reaches 21.7% success rate, outperforming baselines trained on Windows data. The hybrid action mechanism proves critical, reducing error propagation while maintaining execution efficiency. Through this work, we establish a scalable paradigm that effectively bridges primitive low-level GUI interactions and high-level programmatic intelligence, paving the way for more robust, efficient, and unified computer-use agents that can adapt to diverse environments and complex user tasks in the real world.
@@
Figure 1: (a) UltraCUA’s performance on OSWorld and WindowsAgentArena; (b) Comparison between existing GUI Agents and UltraCUA. Pure low-level actions lead to cascade errors, while UltraCUA is faster and stronger.
@@ 1 Introduction
Computer-use automation has emerged as a critical capability for enabling autonomous agents to interact with the vast ecosystem of desktop and web applications that humans use daily (Hong et al., 2023; Shaw et al., 2024; Zhang et al., 2023). However, current computer-use agents (CUAs) face a fundamental limitation: they operate exclusively through primitive actions such as clicking, typing, and scrolling (Rawles et al., 2024; Koh et al., 2024). This constraint creates a significant performance gap compared to agents that leverage rich programmatic interfaces—APIs, MCP servers, and tools—to accomplish complex tasks efficiently (Qin et al., 2023b; Schick et al., 2023b).
@@
The reliance on primitive actions introduces critical challenges. First, lengthy execution chains accumulate errors that cascade into failures—a single misplaced click can derail an entire task (Zheng et al., 2024; Yan et al., 2023). Second, operations that could be accomplished with a single programmatic call require dozens of GUI actions, creating performance bottlenecks. For example, extracting data from multiple spreadsheets requires a traditional CUA to navigate menus, select cells individually, copy values, switch applications, and paste content—each action a potential failure point. In contrast, an agent with spreadsheet APIs could accomplish this reliably with far fewer operations. This efficiency gap is stark: while other agents leveraging programmatic interfaces exceed 80% success on benchmarks like GAIA (Mialon et al., 2024; Zhang et al., 2025), GUI-only computer-use agents remain fundamentally limited, motivating our unified approach that combines GUI generality with programmatic efficiency.
In this paper, we bridge this capability gap through hybrid action, seamlessly integrating GUI primitives with high-level programmatic tool calls. Rather than treating these as mutually exclusive options, our approach enables a strategic combination of both modes. Agents learn to leverage programmatic tool calls when they provide clear efficiency gains, while retaining GUI interactions for universal coverage and fine-grained control. To summarize, our technical contributions include:
@@
  • An automated pipeline for collecting programmatic tools that scales beyond manually curated sets (Qin et al., 2023a; Tang et al., 2023). Our system extracts tools from software documentation, integrates open-sourced implementations, and employs coding agents to generate new tools on demand. This scalable pipeline produces hundreds of tools across diverse environments, from OSWorld’s Ubuntu applications to WindowsAgentArena’s Windows ecosystem.
  • A dual-pipeline synthetic data engine for verifiable computer-use task generation. Large-scale task synthesis for CUA training is challenging due to the complexity of verifying task completion in dynamic environments. To address this, we develop two complementary pipelines producing 16,000+ verified tasks. The first pipeline employs an instruction-first strategy where agents explore computer environments and propose tasks based on observed states, with trajectories verified by evaluator agents. The second pipeline uses an evaluator-first strategy, collecting atomic verification functions (e.g., checking Chrome URLs, verifying file paths, validating image attributes) from environments, then reprogramming (e.g., modifying parameters) and composing (e.g., combining multiple checks) them to create complex evaluation criteria. LLMs generate tasks satisfying these pre-defined evaluators, ensuring reliable trajectory assessment for training.
  • A large-scale hybrid action trajectory collection. Existing computer-use datasets contain only pure GUI action sequences, lacking demonstrations of programmatic tool integration. We collect 20,000+ successful trajectories by combining a powerful planner model (OpenAI o3) with a state-of-the-art grounding model (GTA1-7B (Yang et al., 2025a))—a simple yet effective agentic framework. The planner selects between programmatic tool calls and low-level GUI actions based on task context, while the grounder ensures accurate GUI execution. This dataset enables training models to seamlessly alternate between action modes for optimal task completion.
  • A foundation agent model with hybrid action trained using the programmatic tools, synthetic tasks, and rollout trajectories described above. We train models at two scales (7B and 32B) through supervised fine-tuning on the high-quality trajectories from our collection, followed by online reinforcement learning on our verifiable synthetic tasks. This two-stage approach produces agents that effectively select between GUI primitives and programmatic tool calls based on task context.
Experiments demonstrate substantial improvements over state-of-the-art CUAs. On OSWorld (Xie et al., 2024), our models achieve an average 22% relative improvement over their base models across both scales. Notably, out-of-domain evaluation on WindowsAgentArena (Bonatti et al., 2024)—without any Windows-specific training—shows our 7B model reaches 21.7% success rate, outperforming baselines trained on Windows data. These results validate that hybrid action provides consistent benefits across model scales and platforms. Our code, models, and datasets will be released to facilitate future research.
@@ 2 Methodology
Figure 2: An overview of UltraCUA’s design. The agent adaptively switches between visual grounding and programmatic tool call, establishing the hybrid action mechanism.
@@ 2.3.4 Stage 2: Online Reinforcement Learning
For policy optimization, we employ a variant of GRPO (Shao et al., 2024) inspired by DAPO (Rosset et al., 2024), with key modifications for our hybrid action setting. We remove KL regularization and implement a clip-higher strategy to encourage exploration of diverse action sequences.

### v2

@@ Abstract
Computer-use agents face a fundamental limitation. They rely exclusively on primitive GUI actions (click, type, scroll), creating brittle execution chains prone to cascading failures. While API-driven agents harness rich capabilities through structured interfaces and tools, computer-use agents remain constrained to low-level visual interactions. We present UltraCUA, a foundation model that transcends this limitation through hybrid action—seamlessly unifying primitive GUI operations with high-level tool execution. Our innovation rests on four critical advances. First, an automated pipeline extracts and scales tool capabilities from software documentation and code repositories. Second, a synthetic data engine produces 17,000+ verifiable tasks capturing real-world computer-use complexity. Third, comprehensive hybrid action trajectory collection incorporates both GUI primitives and strategic tool calls. Fourth, a two-stage training methodology combines supervised fine-tuning with online reinforcement learning, enabling intelligent action selection between GUI and API. Evaluation with our 7B and 32B UltraCUA models reveals transformative performance gains. On OSWorld, UltraCUA achieves 22% relative improvement while executing 11% faster than existing approaches, averagely. Cross-domain validation on WindowsAgentArena demonstrates robust generalization with 21.7% success rate, surpassing Windows-trained baselines. The hybrid action paradigm proves essential, reducing error propagation while improving execution efficiency. This work establishes a scalable paradigm bridging primitive GUI interactions and high-level tool intelligence, enabling more resilient and adaptable computer use agents for diverse environments and complex real-world tasks.
@@
Figure 1: (a) UltraCUA demonstrates superior performance on OSWorld and WindowsAgentArena benchmarks; (b) Comparison of execution paradigms between existing GUI agents and UltraCUA. Traditional agents using pure low-level actions suffer from cascading errors, while UltraCUA’s hybrid approach achieves faster and more robust execution.
@@ 1 Introduction
Computer-use automation has emerged as a transformative capability for autonomous agents. It enables interaction with the vast ecosystem of desktop and web applications that define modern digital workflows (Hong et al., 2023; Shaw et al., 2024; Zhang et al., 2023). However, current computer-use agents (CUAs) face a critical architectural limitation. They operate exclusively through primitive visual actions such as clicking, typing, and scrolling (Rawles et al., 2024; Koh et al., 2024). This constraint creates a profound performance chasm compared to API-driven agents that harness rich tool-based interfaces—MCP servers and specialized tools to accomplish complex tasks with efficiency (Qin et al., 2023b; Schick et al., 2023b).
@@
The exclusive reliance on primitive actions introduces fundamental challenges that severely limit agent effectiveness. First, extended execution chains accumulate errors that propagate into catastrophic failures. A single misplaced click can completely derail complex workflows (Zheng et al., 2024; Yan et al., 2023). Second, operations achievable through single tool calls require dozens of fragile GUI interactions. This creates severe performance bottlenecks. For instance, extracting data from multiple spreadsheets forces traditional CUAs to navigate complex menus, manually select individual cells, copy values, switch applications, and paste content. Each step represents a critical failure point. Conversely, an agent with native spreadsheet APIs could accomplish this task reliably with minimal operations. This efficiency disparity is striking. While agents leveraging tool-based interfaces consistently exceed 80% success on challenging benchmarks like GAIA (Mialon et al., 2024; Zhang et al., 2025), GUI-constrained computer-use agents remain fundamentally bottlenecked. This compels our unified approach that marries GUI universality with tool-based precision.
In this paper, we eliminate this capability gap through hybrid action, a paradigm that recognizes the complementary strengths of both execution modalities. Our approach seamlessly integrates GUI primitives with high-level tool execution, unlocking the best of both worlds: the universal applicability of visual interactions and the efficiency of direct tool access. Rather than treating these paradigms as mutually exclusive alternatives, we enable intelligent orchestration that maximizes task success while minimizing execution overhead. Agents learn to strategically deploy tool calls when they offer clear efficiency advantages, while preserving GUI interactions for comprehensive coverage and granular control when tools are unavailable or insufficient.
To summarize, our technical contributions include:
@@
  • An automated pipeline for synthesizing tool capabilities that transcends manual curation limitations (Qin et al., 2023a; Tang et al., 2023). Unlike existing approaches that rely on fixed tool sets, our system dynamically extracts tools from software documentation, integrates open-sourced implementations, and employs coding agents to generate tools on demand. This enables scalable tool acquisition across diverse environments, producing hundreds of tools that form a critical foundation for hybrid action execution.
  • A dual-pipeline synthetic data engine that solves the fundamental challenge of verifiable task generation for computer-use agent training. Traditional approaches fail to reliably assess task completion in dynamic environments, creating unreliable training data. We develop two innovative pipelines that together produce 17,000+ tasks with guaranteed verification. The instruction-first pipeline leverages agent exploration to generate contextually relevant tasks, with dedicated evaluators ensuring execution validity. The evaluator-first pipeline reverses this process by collecting atomic verification functions from target environments, then reprogramming or composing them for complex assessment criteria. Tasks are generated to satisfy these pre-built evaluators, ensuring reliable completion detection. This breakthrough eliminates the verification bottleneck that has limited training computer-use agents at scale.
  • A large-scale hybrid action trajectory collection that bridges the critical gap between pure GUI datasets and tool-augmented execution. Existing computer-use datasets contain only primitive GUI action sequences. They lack demonstrations of strategic tool integration, severely limiting agent learning capabilities. We address this fundamental limitation by collecting 20,000+ successful trajectories using an multi-agent framework. Our approach combines a powerful planner (e.g., OpenAI o3) with a state-of-the-art grounding model (e.g., GTA1-7B (Yang et al., 2025a)). The planner intelligently selects between tool calls and GUI actions based on efficiency considerations. The grounder ensures precise execution of chosen actions. This dataset uniquely captures the decision-making patterns required for seamless action mode transitions, enabling agents to learn when and how to leverage hybrid capabilities.
  • An agentic foundation model with hybrid action intelligence that transcends the limitations of GUI-only training. We develop models at two scales (7B and 32B) through our data and tool ecosystem. Our training employs a strategic two-stage methodology. First, supervised fine-tuning on high-quality hybrid trajectories teaches agents when to leverage tools versus GUI actions. Second, online reinforcement learning on verifiable synthetic tasks refines decision-making through on-policy feedback. This approach produces agents that strategically balance tool efficiency with GUI universality.
Experiments reveal the transformative impact of UltraCUA’s hybrid action across diverse benchmarks and platforms. On OSWorld (Xie et al., 2024), our UltraCUA models deliver a compelling 22% relative improvement over base models at 7B and 32B scales, averagely. In terms of cross-platform generalization, UltraCUA-7B achieves 21.7% success on WindowsAgentArena (Bonatti et al., 2024) without any Windows-specific training, surpassing models explicitly trained on Windows data. This demonstrates that UltraCUA’s hybrid action capabilities fundamentally enhance agent adaptability rather than merely optimizing for specific environments. The consistent gains across model scales and operating systems establish UltraCUA as a new paradigm for robust computer-use automation.
@@ 2 Methodology
Figure 2: An overview of UltraCUA’s design. UltraCUA adaptively switches between visual grounding and tool execution based on task requirements. This establishes the hybrid action mechanism that seamlessly integrates GUI primitives with high-level tool capabilities for enhanced computer-use automation across diverse environments.
@@ 2.3.4 Stage 2: Online Reinforcement Learning
For policy optimization, we employ a variant of GRPO (Shao et al., 2024) inspired by DAPO (Yu et al., 2025), with key modifications for our hybrid action setting. We remove KL regularization and implement a clip-higher strategy to encourage exploration of diverse action sequences.